OpenStudy (anonymous):

NO NEED TO SOLVE JUST determine if its geometric series or geometric sequence or arithmetic seriesA fireplace contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, how many bricks are on the 12th row?

OpenStudy (anonymous):

46,42,.... a =46 d = -4 12th row means n= 12 tn = a+(n-1)d tn =46+(11)*(-4) = 46 - 44 = 2 bricks

jimthompson5910 (jim_thompson5910):

You're decreasing the number of bricks by a set amount each time (in this case 4) So this is an arithmetic sequence.